OpenStudy (anonymous):

The slope of the line passing through the points (2, 7) and (-4, 8) is

OpenStudy (sasogeek):

- (1/6)

OpenStudy (sasogeek):

slope = (y2-y1/x2-x1) =(8-7/-4-2) =(1/-6) =-(1/6)